QUESTION: Show= (y − y0* )(y − y1*) . .(y − yn* ) = 5 \frac{5^{n+1}}{2^{n}}*T_{n+1}(x)), x=\frac{y}{5}

it is Part 1 at the bottom
Answer #1

Recull at and are given by X; = cos(2 + 1 1 1 1 +(x) I Itar Solution for given Probleme - -Given that the information thet ThThe, (u) - 64-bit) 10-24). -- (1-0 from goin=0 Ti() = (x-mont) 40 x to an Ao-1 0.0 : زر .از 215 - COD) T2 (W2 = cos la costa)3 como cos (30)= -Coupo sun2an 4.no 1²-31 = (x-v) In-o) (x+) 42 = n(n² 2 / AL (4-4)42 A2 = 4 So, proceding this we have Anal?